OpenStudy (shamil98):

series stuff Determine whether the series is absolutely convergent, conditionally convergent, or divergent. \[\sum_{n=1}^{\infty} ( 1 + \frac{ 1 }{ n })^{n^2} \] Using the DCT i can compare this to \[\lim_{n \rightarrow \infty} (1 + \frac{ 1 }{ n })^n\] but i'm not sure how to find the limit of this one ( i know it converges to e, but how so?)

ganeshie8 (ganeshie8):

isn't it simply the definition of e ?

ganeshie8 (ganeshie8):

By definition : \[\lim_{n \rightarrow \infty} (1 + \frac{ x }{ n })^n=e^x\]

OpenStudy (shamil98):

how do i prove that though?

ganeshie8 (ganeshie8):

you don't prove definitions

OpenStudy (shamil98):

thanks

ganeshie8 (ganeshie8):

Np :) I think you're using "limit test", not DCT